# Vision Transformers on the Edge: A Comprehensive Survey of Model Compression and Acceleration Strategies

### v1

@@ 4.1 Accelerating Non-linear operations
ViT models in CV can mainly be split into two types of operations: linear and non-linear. Optimizing non-linear operations in quantized ViT is as crucial as optimizing linear operations. While low-bit computing units significantly reduce computational complexity and memory footprint, They are primarily designed for linear operations such as matrix multiplications and convolutions. However, non-linear functions, including softmax, GELU, and LayerNorm, are the essential components of ViT architectures yet unexplored largely in these low-bit computing environments. The lack of support for non-linear operations on hardware creates computational bottlenecks during the edge deployment, as non-linear functions often require FP32 operations. This results in latency and increases power and energy consumption, ultimately lessening the quantization’s full benefits on edge deployment. Moreover, during inference on quantized ViT models on edge devices, frequent quantization and dequantization operations surrounding non-linear layers add further inefficiencies, slowing inference and reducing throughput [115]. Researchers developed integer-based approximations for non-linear operations to solve these issues, eliminating the need for frequent FP32 computations in those layers.
@@
As illustrated in Table 11, we introduce integer-only approximations for different hardware platforms to enhance ViT inference efficiency. As we discussed details in section 2.3, FQ-ViT [14] utilized LIS for integer-only variants of the softmax that approximate the exponential component using second-order polynomial coupled with $\log 2$ quantization. For LayerNorm, the authors applied PTF to shift the quantized activations and later computed mean and variance using integer arithmetic. However, this approach seems hardware efficient; their methods of practical deployment on edge platforms are unclear. Additionally, I-ViT [87] calculates the square root in LayerNorm using an integer-based iterative method. The authors then introduced ShiftGELU, which used sigmoid-based approximations for GELU approximations. I-ViT utilized NVIDIA RTX 2080 Ti GPU as a hardware platform to evaluate their method.
@@
PackQViT [106] is an extended version of the FQ-ViT concept that also leverages second-order polynomial approximations more straightforwardly, replacing the usual constant $e$ with 2 in the softmax. Although PackQViT requires training, the simplification ensures no accuracy loss. EdgeKernel [116] addresses precision challenges in softmax computations by optimizing the selection of the bit shift parameter on Apple A13 and M1 chips, ensuring high accuracy while minimizing significant bit truncations. Additionally, it employs asymmetric quantization for LayerNorm inputs, converting them to a uint16 format to enhance computational efficiency while maintaining data integrity.
@@
SOLE [117] and SwiftTron [118] both are utilized application-specific integrated circuit (ASIC) platforms. SOLE optimizes the software perspectives, introducing E2Softmax with $\log 2$ quantization to avoid traditional FP32 precision in softmax layers. Additionally, SOLE designed a two-stage LayerNorm unit using PTF factors. However, SwiftTron focused on designing customized hardware for ASIC to efficiently use non-linear operations, even in FP32, which accounts for diverse scaling factors in performing correct computations.
@@ 4.2 Current Accelerating Techniques on ViT
Table 12: The overview of current accelerating techniques for ViT on edge devices. In the baseline models, B denotes Base; S denotes Small, and T denotes Tiny versions. Five types of hardware devices are used in existing hardware-software co-design frameworks: GPU ($\spadesuit$), EdgeGPU ($\blacktriangle$), CPU ($\bigstar$), FPGA ($\clubsuit$), and AMD Versal Adaptive Compute Acceleration (ACAP) ($\blacklozenge$).
@@
4.2.1 Software Optimization in Software-Hardware Co-design
@@
Quantization-Based Software Acceleration VAQF [119] utilized binary quantization for weights and low-precision for the activations in the software part. VAQF automatically outputs the efficient quantization parameters based on the model structure and the expected frame per second (FPS) to meet the hardware specifications. The primary purpose of this work was to achieve high throughput on hardware while maintaining model accuracy. Moreover, Li et al. [120] used a mixed-scheme (fixed+PoT) ViT quantization algorithm that can fully leverage heterogeneous FPGA resources for a target FPS. Both frameworks utilized targeted FPS as their input to achieve maximum hardware efficiency during inference. Additionally, both VAQF and AutoViT-Acc [120] utilized PTQ methods as quantization. EQ-ViT [121] combined latency and accuracy requirements to decide the final quantization strategy leveraging the QAT approach. Additionally, EQ-ViT implemented a profiling-based execution scheduler that dynamically allocates workloads across hardware accelerators.
Pruning-Based Software Acceleration ViT’s self-attention has quadratic complexity concerning input sequence length, leading to high memory bandwidth consumption. Sparse attention mechanisms aim to reduce redundant computation by focusing on pruning. For instance, ViTCoD [23] efficiently employed structured pruned and polarized the attention maps to remove redundant attention scores, creating a more memory-efficient execution. Moreover, HeatViT [20] utilized image-adaptive token pruning, identifying and removing unimportant tokens before transformer blocks using a multi-token selector, dynamically reducing computational complexity. The proposed method is highly inspired by SP-ViT  [38] and uses a similar token selector like SP-ViT.
@@
4.2.2 Hardware Optimization in Software-Hardware Co-design
Hardware optimization plays a crucial role in the software-hardware co-design of ViTs, ensuring efficient execution across different accelerators. Key techniques include C++ based hardware descriptions, high-level synthesis (HLS), and accelerator bitstream generation [119, 120, 20]. Additionally, frameworks leverage AI engine (AIE) kernels [121] and custom hardware units (e.g., SOLE) to optimize execution for edge deployment.
FPGA-Based Hardware Acceleration FPGA is the pioneer for hardware accelerating strategies because of its reconfigurable characteristics. VAQF [119] adopted the quantization schemes from the software part and applied them to the accelerator on the hardware side. Figure 6 illustrates the overview of the VAQF framework. The accelerator’s C++ description was synthesized using the Vivado HLS tool. Initial accelerator parameters focused on maximizing parallelism, but there were adjustments due to Vivado’s placement or routing problems. Successful implementations produced a bitstream file for FPGA deployment. Moreover, as illustrated in Figure 6, Auto-ViT Acc [120] first used the "FPGA Resource Utilization Modeling" module to give performance analysis and calculate the FPS of the FPGA ViT accelerator in software part. Inspired by VAQF, the authors implemented the FPGA accelerator using a C++ hardware description, synthesized through Vitis HLS to generate the final accelerator bitstream.
@@ 4.2.2 Hardware Optimization in Software-Hardware Co-design
Optimizing Non-Linear Operations with Custom Hardware Units Recent studies, such as EQ-ViT, separated matrix multiply (MM) and non-MM by efficiently mapping batch MM (BMM) and convolutions to AIE vector cores. Memory-bounded and non-linear operations are executed within the FPGA’s programmable element (PE). Additionally, EQ-ViT leveraged fine-grained pipeline execution to overlap computation with memory transfers, maximizing resource utilization. A key strength of this framework lies in its hardware mapping methodology, where execution is formulated as a mixed-integer programming (MIP) optimization problem, ensuring that latency and resource constraints are satisfied while maximizing throughput. Likewise, SOLE [121] designed separate units for the proposed E2Softmax and AILayerNorm to perform non-linear operations efficiently. The E2Softmax unit included Log2Exp and an approximate Log-based divider that is implemented in a LUT-free and multiplication-free manner. Additionally, the AILayerNorm unit operates in two stages: the first stage performs statistical calculations, while the second stage applies the affine transformation. Similar to M3ViT [19] also utilized ping pong buffer to pipeline the AILayerNorm unit.
Pure Hardware Accelerators Although most ViT acceleration techniques rely on SW-HW co-design—where software optimization plays a crucial role in achieving high efficiency and accuracy—the potential of pure hardware optimization has gained attention in recent studies. A recent study named ViA [21] addressed issues during data and computations flow through the layers in ViT. The authors utilized a partitioning strategy to reduce the impact of data locality in the image and enhance the efficiency of computation and memory access. Additionally, by examining the computing flow of the ViT, the authors also utilized the half-layer mapping and throughput analysis to lessen the effects of path dependency due to the shortcut mechanism and to maximize the use of hardware resources for efficient transformer execution. The study developed two reuse processing engines with an internal stream, distinguishing them from previous overlaps or stream design patterns drawing from the optimization strategies.
Moreover, ViTA [22] used two sets of MAC units to minimize the off-chip memory accesses. The first set of MAC units representing the hidden layers was broadcasted to the second set of MAC units through a non-linear activation function. That broadcasting approach helped to compute the partial products corresponding to the output layer. The authors allocated these resources to maintain the pipeline technique as if the hidden layer value computations and the output layer partial product computation took equal time. This approach enabled the integration of several mainstream ViT models by only adjusting the configuration.
@@ 4.2.3 Other Techniques
ED-ViT [122] utilized distributed workloads approaches to deploy the ViT mode utilizing multiple tiny edge devices such as Raspberry Pi-4B. The authors partitioned the model into multiple submodels, mapping each submodel to a separate edge device. This distributed execution strategy enables ViTs to achieve efficiency comparable to single powerful edge accelerators like EdgeGPUs or FPGAs while leveraging cost-effective and scalable edge computing resources. Similarly, COSA Plus, proposed by Wang et al. [123], capitalized on high inherent parallelism within ViT models by implementing a runtime-configurable hybrid dataflow strategy. This method dynamically switches between weight-stationary and output-stationary dataflows in a systolic array, optimizing the computational efficiency for matrix multiplications within the attention mechanism. COSA Plus enhances processing element (PE) utilization by adapting data movement patterns to workload variations.
@@ 4.2.4 Discussion
VAQF and Auto-ViTAcc apply PTQ optimization to reduce precision while maintaining accuracy, while HeatViT and EQ-ViT optimize workload distribution through kernel-level techniques and adaptive token pruning. Additionally, the quantization-based acceleration has shown significant improvements in energy efficiency. However, MoE introduces irregular memory access patterns, leading to suboptimal hardware utilization. While ping-pong buffering is a well-established technique to mitigate memory stalls, further optimizations are required to exploit hardware parallelism and memory efficiency.
@@
Most current acceleration techniques for ViTs are highly customized for specialized hardware, such as FPGAs and ASICs, limiting their deployment flexibility. Table 13 highlights the trade-offs between accuracy, adaptability, and hardware compatibility in SW-HW co-design techniques for ViTs. However, few approaches support cross-platform compatibility, making their extension to diverse edge devices challenging. Among the surveyed techniques, PIVOT [124] emerges as the most flexible solution, as it maintains high accuracy while being compatible with GPPs, unlike ViTCoD, HeatViT, and VAQF, which are hardware-specialized acceleration techniques. Additionally, Table 13 highlights how different acceleration techniques prioritize computations, offering insights into their adaptability across various deployment scenarios.
@@ 4.3 Performance Analysis for Accelerating Techniques
Framework Device Precision Frequency (MHZ) Resource Utilization FPS Power (W) Energy Efficiency (FPS/W) Throughput (GOPs) Speedup ($\uparrow$)
@@ 4.3.1 Resource Utilization
From the observation of Table 14, HeatViT (2066 DSPs) [20] and Zhang et al. (2147 DSPs) [125] demonstrate the highest DSP consumption, indicating their reliance on intensive parallel processing to accelerate transformer computations although both used different FPGA variants. In terms of on-chip memory usage, ViA (1002 BRAMs) [21] and VAQF (565.5 BRAMs) [119] exhibit significant BRAM consumption, emphasizing a design strategy that prioritizes data locality to minimize off-chip memory access latency. In contrast, EQ-ViT (16 BRAMs) [121] utilizes remarkably low BRAM, likely due to its two-level optimization kernels—leveraging both single artificial intelligence engines (AIEs) and AIE array levels. Additionally, Auto-ViT-Acc [120] (186 KLUTs) and HeatViT [20] (161.4 KLUTs) indicate that they need to perform significant logical operations to implement their mixed precision quantization and token pruning to deploy on edge.
@@ 4.3.2 Energy Efficiency
Table 14 also indicates the energy efficiency of the ViT accelerating techniques. In the current studies, energy efficiency was measured in two ways: using throughput (GOP/J) and using FPS (FPS/watt). We also include the power consumption of the accelerating techniques for better accountability. It is perhaps difficult to conclude about energy efficiency when different edge targets are used for the CV task. From Table 14, VAQF [119], Auto-ViT-Acc [120], and HeatViT [20] utilized the same FPGA AMD ZCU102 board with the same number of resources and frequency. We can observe that VAQF outperforms Auto-ViT-Acc and HeatViT regarding energy and power usage. However, we were unable to find any energy comparison from the original paper for ASIC-based accelerating techniques (e.g., ViTCoD, SOLE).

### v2

@@ 3.3 Common Optimization Techniques
Although software frameworks enable optimized ViT execution with hardware-aware features, achieving real-time performance on edge devices often requires low-level, hardware-specific acceleration techniques. Section 4 explores dedicated acceleration methods for deploying ViTs on edge platforms.
@@ 4.1 Accelerating Non-linear operations
ViT models in CV can mainly be split into two types of operations: linear and non-linear. Optimizing non-linear operations in quantized ViT is as crucial as optimizing linear operations. While low-bit computing units significantly reduce computational complexity and memory footprint, they are primarily designed for linear operations such as matrix multiplications and convolutions. However, non-linear functions, including softmax, GELU, and LayerNorm, are the essential components of ViT architectures yet largely unexplored in these low-bit computing environments. The lack of support for non-linear operations on hardware creates computational bottlenecks during the edge deployment, as non-linear functions often require FP32 operations. This results in latency and increases power and energy consumption, ultimately lessening the full benefits of quantization on edge deployment. Moreover, during inference on quantized ViT models on edge devices, frequent quantization and dequantization operations surrounding non-linear layers add further inefficiencies, slowing inference and reducing throughput [118]. Researchers developed integer-based approximations for non-linear operations to solve these issues, eliminating the need for frequent FP32 computations in those layers.
@@
As illustrated in Table 14, we introduce integer-only approximations for different hardware platforms to enhance ViT inference efficiency. As we discussed details in section 2.3, FQ-ViT [14] utilized LIS for integer-only variants of the softmax that approximate the exponential component using a second-order polynomial coupled with $\log 2$ quantization. For LayerNorm, the authors applied PTF to shift the quantized activations and later computed mean and variance using integer arithmetic. However, this approach seems hardware efficient; their methods of practical deployment on edge platforms are unclear. Additionally, I-ViT [93] calculates the square root in LayerNorm using an integer-based iterative method. The authors then introduced ShiftGELU, which used sigmoid-based approximations for GELU approximations. I-ViT utilized NVIDIA RTX 2080 Ti GPU as a hardware platform to evaluate their method.
@@
PackQViT [109] is an extended version of the FQ-ViT concept that also leverages second-order polynomial approximations more straightforwardly, replacing the usual constant $e$ with 2 in the softmax. Although PackQViT requires training, the simplification ensures no accuracy loss. EdgeKernel [119] addressed precision challenges in softmax computations by optimizing the selection of the bit shift parameter on Apple A13 and M1 chips, ensuring high accuracy while minimizing significant bit truncations. Additionally, it employs asymmetric quantization for LayerNorm inputs, converting them to a uint16 format to enhance computational efficiency while maintaining data integrity.
@@
SOLE [120] and SwiftTron [121] both are utilized application-specific integrated circuit (ASIC) platforms. SOLE optimizes the software perspectives, introducing E2Softmax with $\log_{2}$ quantization to avoid traditional FP32 precision in softmax layers. Additionally, SOLE designed a two-stage LayerNorm unit using PTF factors. However, SwiftTron focused on developing customized hardware for ASIC to efficiently use non-linear operations, even in FP32, which accounts for diverse scaling factors in performing correct computations.
@@ 4.2 Current Accelerating Techniques on ViT
Table 15: The overview of current accelerating techniques for ViT on edge devices. In the baseline models, B denotes Base; S denotes Small, and T denotes Tiny versions. Five types of hardware devices are used in existing SW-HW co-design frameworks: GPU ($\spadesuit$), EdgeGPU ($\blacktriangle$), CPU ($\bigstar$), FPGA ($\clubsuit$), and AMD Versal Adaptive Compute Acceleration (ACAP) ($\blacklozenge$).
@@
4.2.1 Software Optimization in SW-HW Co-design
@@
Quantization-Based Software Acceleration VAQF [122] utilized binary quantization for weights and low-precision for the activations in the software part. VAQF automatically outputs the efficient quantization parameters based on the model structure and the expected frame per second (FPS) to meet the hardware specifications. The primary purpose of this work was to achieve high throughput on hardware while maintaining model accuracy. Moreover, Li et al. [123] used a mixed-scheme (fixed+PoT) ViT quantization algorithm that can fully leverage heterogeneous FPGA resources for a target FPS. Both frameworks utilized targeted FPS as their input to achieve maximum hardware efficiency during inference. Additionally, both VAQF and AutoViT-Acc [123] utilized PTQ methods as quantization. EQ-ViT [124] combined latency and accuracy requirements to decide the final quantization strategy , leveraging the QAT approach. Additionally, EQ-ViT implemented a profiling-based execution scheduler that dynamically allocates workloads across hardware accelerators.
Pruning-Based Software Acceleration ViT’s self-attention has quadratic complexity concerning input sequence length, leading to high memory bandwidth consumption. Sparse attention mechanisms aim to reduce redundant computation by focusing on pruning. For instance, ViTCoD [23] efficiently employed structured pruning and polarized the attention maps to remove redundant attention scores, creating a more memory-efficient execution. Moreover, HeatViT [20] utilized image-adaptive token pruning, identifying and removing unimportant tokens before transformer blocks using a multi-token selector, dynamically reducing computational complexity. The proposed method is highly inspired by SP-ViT  [37] and uses a similar token selector like SP-ViT.
@@
4.2.2 Hardware Optimization in SW-HW Co-design
Hardware optimization plays a crucial role in the SW-HW co-design of ViTs, ensuring efficient execution across different accelerators. Key techniques include C++ based hardware descriptions, high-level synthesis (HLS), and accelerator bitstream generation [122, 123, 20]. Additionally, frameworks leverage AI engine (AIE) kernels [124] and custom hardware units (e.g., SOLE) to optimize execution for edge deployment.
ASIC-Based Hardware Acceleration ASIC-based accelerators have become a cornerstone in deploying deep learning models on edge devices, with the majority of advancements driven by leading technology companies. Notable examples include Google’s TPU [127], IBM’s TrueNorth [127], and NVIDIA’s open-source NVDLA [128]. Academic research has also contributed with efficient designs such as Origami [129] and ZASCA [130], which are tailored for low-power convolutional inference.
Google’s TPUs have released several versions [131]. TPUv1, released in 2016, featured a 256$\times$256 systolic array of 8-bit fixed-point multipliers, delivering up to 92 TOPS at 75 W—well-suited for accelerating matrix-heavy operations like those found in ViT inference. TPUv2 and TPUv3 introduced support for FP operations and higher throughput. For resource-constrained environments, Google developed the EdgeTPU, a lightweight variant tailored for edge AI workloads.
As ViTs grow in popularity, their unique architectural demands have motivated the development of ViT-specific ASIC accelerators. Unlike CNN-based accelerators, ViT inference requires efficient handling of attention mechanisms and normalization layers. Recent efforts such as ViTCoD [23] and SOLE[120] addressed these challenges through SW-HW co-design strategies and have demonstrated promising ASIC implementations for ViT acceleration at the edge.
FPGA-Based Hardware Acceleration FPGA is the pioneer for hardware accelerating strategies because of its reconfigurable characteristics. VAQF [122] adopted the quantization schemes from the software part and applied them to the accelerator on the hardware side. Figure 12 illustrates the overview of the VAQF framework. The accelerator’s C++ description was synthesized using the Vivado HLS tool. Initial accelerator parameters focused on maximizing parallelism, but there were adjustments due to Vivado’s placement or routing problems. Successful implementations produced a bitstream file for FPGA deployment. Moreover, as illustrated in Figure 12, Auto-ViT Acc [123] first used the "FPGA Resource Utilization Modeling" module to give performance analysis and calculate the FPS of the FPGA ViT accelerator in the software part. Inspired by VAQF, the authors implemented the FPGA accelerator using a C++ hardware description, synthesized through Vitis HLS to generate the final accelerator bitstream.
@@ 4.2.2 Hardware Optimization in SW-HW Co-design
Optimizing Non-Linear Operations with Custom Hardware Units Recent studies, such as EQ-ViT, separated matrix multiply (MM) and non-MM by efficiently mapping batch MM (BMM) and convolutions to AIE vector cores. Memory-bounded and non-linear operations are executed within the FPGA’s programmable element. Additionally, EQ-ViT leveraged fine-grained pipeline execution to overlap computation with memory transfers, maximizing resource utilization. A key strength of this framework lies in its hardware mapping methodology, where execution is formulated as a mixed-integer programming (MIP) optimization problem, ensuring that latency and resource constraints are satisfied while maximizing throughput. Likewise, SOLE [124] designed separate units for the proposed E2Softmax and AILayerNorm to perform non-linear operations efficiently. The E2Softmax unit included Log2Exp and an approximate Log-based divider that is implemented in a LUT-free and multiplication-free manner. Additionally, the AILayerNorm unit operates in two stages: the first stage performs statistical calculations, while the second stage applies the affine transformation. Similar to M3ViT [19], SOLE also utilized ping pong buffer to pipeline the AILayerNorm unit.
Pure Hardware Accelerators Although most ViT acceleration techniques rely on SW-HW co-design—where software optimization plays a crucial role in achieving high efficiency and accuracy—the potential of pure hardware optimization has gained attention in recent studies. A recent study named ViA [21] addressed issues during data flow through the layers in ViT. The authors utilized a partitioning strategy to reduce the impact of data locality in the image and enhance the efficiency of computation and memory access. Additionally, by examining the computing flow of the ViT, the authors also utilized the half-layer mapping and throughput analysis to lessen the effects of path dependency due to the shortcut mechanism and to maximize the use of hardware resources for efficient transformer execution. The study developed two reuse processing engines with an internal stream, distinguishing them from previous overlaps or stream design patterns from the optimization strategies.
Moreover, ViTA [22] used two sets of MAC units to minimize the off-chip memory accesses. The first set of MAC units representing the hidden layers was broadcast to the second set of MAC units through a non-linear activation function. That broadcasting approach helped to compute the partial products corresponding to the output layer. The authors allocated these resources to maintain the pipeline technique as if the hidden layer value computations and the output layer partial product computation took equal time. This approach enabled the integration of several mainstream ViT models by only adjusting the configuration.
@@ 4.2.3 Other Techniques
ED-ViT [125] utilized distributed workloads approaches to deploy the ViT model utilizing multiple tiny edge devices such as Raspberry Pi-4B. The authors partitioned the model into multiple submodels, mapping each submodel to a separate edge device. This distributed execution strategy enables ViTs to achieve efficiency comparable to single powerful edge accelerators like EdgeGPUs or FPGAs while leveraging cost-effective and scalable edge computing resources. Similarly, COSA Plus, proposed by Wang et al. [126], capitalized on high inherent parallelism within ViT models by implementing a runtime-configurable hybrid dataflow strategy. This method dynamically switches between weight-stationary and output-stationary dataflows in a systolic array, optimizing the computational efficiency for matrix multiplications within the attention mechanism. COSA Plus enhances processing element (PE) utilization by adapting data movement patterns to workload variations.
@@ 4.2.4 Discussion
VAQF and Auto-ViTAcc apply PTQ optimization to reduce precision while maintaining accuracy, while HeatViT and EQ-ViT optimize workload distribution through kernel-level techniques and adaptive token pruning. On the other hand, MoE introduces irregular memory access patterns, leading to suboptimal hardware utilization. While ping-pong buffering is a well-established technique to mitigate memory stalls, further optimizations are required to exploit hardware parallelism and memory efficiency.
@@
Most current acceleration techniques for ViTs are highly customized for specialized hardware, such as FPGAs and ASICs, limiting their deployment flexibility to other platforms. Table 16 highlights the trade-offs between accuracy, adaptability, and hardware compatibility in SW-HW co-design techniques for ViTs. However, limited support for cross-platform compatibility poses challenges for extending these methods to a wide range of edge devices. Among the surveyed techniques, PIVOT [132] emerges as the most flexible solution, as it maintains high accuracy while being compatible with GPPs, unlike ViTCoD, HeatViT, and VAQF, which are hardware-specialized acceleration techniques. Additionally, Table 16 highlights how different acceleration techniques prioritize computations, offering insights into their adaptability across various deployment scenarios.
@@ 4.3 Performance Analysis for Accelerating Techniques
Framework Device Precision Frequency (MHZ) Resource Utilization FPS Power (W) Energy (FPS/W) Throughput (GOPs) Speedup ($\uparrow$)
@@ 4.3.1 Resource Utilization
From the observation of Table 17, HeatViT (2066 DSPs) [20] and Zhang et al. (2147 DSPs) [133] demonstrate the highest DSP utlization, indicating their reliance on intensive parallel processing to accelerate transformer computations, although both used different FPGA variants. In terms of on-chip memory usage, ViA (1002 BRAMs) [21] and VAQF (565.5 BRAMs) [122] exhibit significant BRAM consumption, emphasizing a design strategy that prioritizes data locality to minimize off-chip memory access latency. In contrast, EQ-ViT (16 BRAMs) [124] utilizes remarkably low BRAM, likely due to its two-level optimization kernels, leveraging both single artificial intelligence engines (AIEs) and AIE array levels. Additionally, Auto-ViT-Acc [123] (186 KLUTs) and HeatViT [20] (161.4 KLUTs) indicate that they need to perform significant logical operations to implement their mixed precision quantization and token pruning to deploy on edge.
@@ 4.3.2 Energy Efficiency
Table 17 also indicates the energy efficiency of the ViT accelerating techniques. In the current studies, energy efficiency was measured in two ways: using throughput (GOP/J) and using FPS (FPS/watt). We also include the power consumption of the accelerating techniques for better accountability. It is perhaps difficult to conclude about energy efficiency when different edge targets are used for the CV task. As illustrated in Table 17, VAQF [122], Auto-ViT-Acc [123], and HeatViT [20] utilized the same FPGA AMD ZCU102 board with the same number of resources and frequency. We can observe that VAQF outperforms Auto-ViT-Acc and HeatViT in terms of energy and power usage. However, we were unable to find any energy comparison from the original paper for ASIC-based accelerating techniques (e.g., ViTCoD, SOLE).